A total of $10,000 is invested in two funds paying 5% and 7% annual interest. The combined annual interest is $644. How much of the $10,000 is invested in each fund?
Answer by checkley71(8403)

.07X+.05(10000-X)=644
.07X+500-.05X=644
.02X=644-500
.02X=144
X=144/.02
X=7200 ANSWER FOR THE AMOUNT INVESTED @ 7%.
10000-7200=2800 INVESTED @ 5%.
PROOF.
.07*7200+.05*2800=644
504+140=644
644=644